solve the equation by completing the square.
x(square) + x - 3/4= 0

Answer by venugopalramana(3286)

solve the equation by completing the square.
x(square) + x - 3/4= 0......multiplying with 4
4x^2+4x-3=0
{(2x)^2+2*(2x)+1^2}-1-3=0
(2x+1)^2=4
2x+1=+2....or....-2
2x=2-1=1.....or......-2-1=-3
x=1/2......or......-3/2